# Hallucination, reliability, and the role of generative AI in science

### v1

@@ 7 Error mitigation and the construction of reliable science
Generative AI is increasingly central to scientific inquiry, yet the spectre of hallucination has raised legitimate doubts about its reliability. I have argued that while corrosive hallucination presents a novel epistemic threat, it is not a reason to adopt a wholesale form of skepticism about the use of generative AI models in science. Although hallucinations complicate the task of shoring up our inferences, scientific workflows have been developed to meet the challenge. As the cases of AlphaFold and GenCast illustrate, these methods do not eliminate error, but they do ensure that errors remain managable.
@@
Although these strategies do not rely on aligning individual model parameters with interpretable features of the world, they are nevertheless more sophisticated than the kind of brute inductivism that Duede has criticized. The mechanisms by which these systems manage error are not ad hoc. They draw on independently supported theoretical knowledge about the target domain, and integrate that knowledge into scientific workflow at during model development and in the interpretation of results.
The view developed here is largely aligned with the what Dúran and Formanek call computational reliabilism, a view about scientific epistemology that was initially developed in order to explore what justification we can have for believing the results of large, epistemically opaque computer simulations. But at least some epistemically opaque computer simulations have what Creel [2020] calls functional transparency. We know which algorithm they are designed to implement; the challenge is knowing whether they implement it correctly. When working with DNN-based models, the problem is worse, because DNNs are not functionally transparent. We have no compact, cognitively accessible representation of the algorithm they implement. This fact has led many scientists and philosophers to worry that AI-driven science is a kind of blind data-driven procedure where prediction is all that matters, and theory holds no sway. My arguments concerning the problem of hallucination is another way of showing, as Andrews [2023] compellingly argues, AI does not offer a theory-free route to knowledge. On the contrary, generative models gain scientific legitimacy precisely when they are disciplined by theory.
@@
Given the rapid pace of AI development, there is a genuine risk that the methodological norms that guide scientific modeling today will not apply to the science of tomorrow. It is important, therefore, to qualify any strong conclusions. In that spirit, I want to flag a position that may seem naturally aligned with the argument of this paper, but which I do not endorse. It is the view that only narrow, domain-specific models can serve as reliable sources of scientific knowledge. While domain-specific models currently outperform general-purpose models in this respect—largely because it is easier to construct error-mitigating workflows around them—I remain agnostic about whether similar forms of reliability might eventually be achieved with general-purpose architectures. There are reasons to take this possibility seriously. Multi-modal models such as GPT-4 and Gemini 2.5 have so-called ’emergent’ capacities that cannot be replicated by chaining together domain-specific tools. Their integration of language, image, and video appears to produce synergies within the model’s latent space that surpass what modular composition can deliver. Whether chemical representations could be treated as an additional “modality” in this sense—and whether that integration would yield genuinely new epistemic benefits—remains an open and intriguing question.
Another way that general-purpose models may become reliable contributors to scientific inference is by means of something like the competitive dynamics among human scientists which, when things go well, give rise to an error-correcting mechanism on a social scale. The Virtual Labs framework [Swanson et al., 2024] exemplifies the possibility of implementing something like this in AI agents: a system of LLMs that pose hypotheses, subject them to criticism by other LLMs, refine the hypotheses, and then eventually farm them out to domain-specific tools like Rosetta and AlphaFold to make predictions, which are then tested by human experimentalists. Though preliminary and not yet peer-reviewed, such systems hint at novel epistemic architectures—ones in which general-purpose models play a meaningful role within a cutting-edge scientific discovery. These developments are too recent to warrant strong conclusions, but they caution against assuming that narrowness is the only viable path.
This paper is intended as a partial answer to skeptical concerns about the use of generative AI in science. Nevertheless, I would like to add a cautionary note of my own, which was inspired by discussions at a recent meeting concerned with European science funding. While enthusiasm for AI-driven science grows, there is a risk of reallocating funding away from traditional theory-building and experimentation. Yet, as this paper has emphasized, generative AI models achieve reliability precisely because they are embedded in theoretical frameworks and validated by empirical experimentation. Neglecting these foundational methods would be unwise, because it would weaken the epistemic scaffolding upon which AI-based inference depends.
@@
The central argument of this paper reflects the Deweyan insight quoted in the epigraph: scientific knowledge is not a matter of grasping antecedently given certainties, but of developing ways to make sure—methods for identifying error and managing uncertainty. Generative AI calls for new methods of doing that work, and simultaneously makes that work more difficult. Nevertheless, as AlphaFold, GenCast, and emerging systems may yet show, there is reason to hope that we will be equal to the task.

### v2

@@ 7 Error mitigation and the construction of reliable science
Generative AI is increasingly central to scientific inquiry, yet the specter of hallucination has raised legitimate doubts about its reliability. I have argued that while hallucination presents a novel epistemic threat, it is not a reason to adopt wholesale skepticism about the use of generative AI in science. The challenge can be addressed by shifting from a data-centric to a phenomenon-centric conception of hallucination. Rather than assessing model outputs by their deviation from training data, we must evaluate them by their correspondence with target phenomena. This shift opens space for alternative reliability strategies, such as theory-guided training and confidence-based screening, that establish systematic output-to-target connections despite opacity. As the cases of AlphaFold and GenCast demonstrate, these methods do not eliminate error, but they make errors anticipatable and manageable.
@@
Although these strategies do not rely on aligning individual model parameters with interpretable features of the world, they are nevertheless more sophisticated than the kind of brute inductivism that Duede has criticized. The mechanisms by which these systems manage error are not ad hoc. They draw on independently supported theoretical knowledge about target phenomena and integrate that knowledge into scientific workflows during model development and in the interpretation of results.
Earlier I invoked Durán’s computational reliabilism to explain how AlphaFold achieves reliability through workflow design despite opacity. Durán and Formanek (2018) originally developed this framework for opaque computer simulations, and Durán [2025] has extended it to algorithms more broadly. Computational reliabilism emphasizes that reliability assessment should focus on the process rather than the algorithm in isolation, where “process” encompasses the broader socio-techno-scientific context in which algorithms are designed, used, and maintained.
My contribution extends this framework to address hallucinations in generative AI for science. The distinctive challenge is that hallucinations must be assessed by comparing outputs to target phenomena rather than training data. This comparison is operationalizable only where we possess sophisticated theoretical knowledge about the target system. Knowing that proteins fold according to thermodynamic principles, that weather obeys conservation laws, or that molecular bonds have characteristic geometries is what enables the two reliability strategies examined here. Theory-guided training works by embedding constraints derived from our understanding of target phenomena into the model’s architecture and loss functions. Confidence-based screening works because we can calibrate uncertainty metrics against theoretical expectations about when predictions should and shouldn’t be reliable. Without mature domain theory, we lack the resources to establish systematic connections between model outputs and target phenomena.
This explains why these strategies succeed in structural biology and meteorology but cannot be assumed to work universally. Both domains benefit from decades of theoretical development: protein folding thermodynamics, structural biochemistry, atmospheric physics, fluid dynamics. This theoretical maturity provides the scaffolding needed to discipline generative models. Where such theory is absent or immature, comparable reliability may be unattainable regardless of computational resources or data availability. Understanding reliability requires assessing the workflow as a whole, not just the model in isolation. Consider Isomorphic Labs, one of the most prominent companies building drug discovery pipelines around AlphaFold-style models. In a recent interview, Rebecca Paul, the company’s head of medicinal chemistry, explained that AlphaFold predictions with binding probability scores below 0.7 are systematically filtered out before any synthesis occurs [Paul and Jaderberg, 2025]. Predictions above this threshold are then validated experimentally through X-ray crystallography, with results feeding back into model refinement. The model output is just one step in a larger process that includes filtering based on confidence thresholds, experimental validation, and iterative improvement. Reliable inference emerges from this extended workflow, not from the model’s computation alone.
This observation connects to longstanding debates in reliabilist epistemology about process individuation [Goldman, 1986]. What counts as “the process” that produces belief? Processes can be individuated more narrowly (the model’s computation) or more broadly (the theory-informed workflow connecting outputs to phenomena). In traditional scientific contexts, we assess experimental procedures as wholes, not individual instruments in isolation. The same principle applies here: the appropriate unit of assessment is the theory-informed, uncertainty-aware, iteratively validated workflow—not the model’s internal mechanisms. This is why opacity, though epistemically significant, does not preclude reliable inference. The workflow, infused with theoretical knowledge at multiple points, is what must be reliable.
These conclusions about workflows and process individuation apply clearly to the current generation of generative AI systems. But given how rapidly AI architectures are evolving, I want to make two additional claims about the intended scope of the arguments above.
First, about the workflows: The strategies examined here (theory-guided training and confidence-based screening) depend fundamentally on sophisticated theoretical knowledge about target phenomena. Establishing reliable output-to-target connections requires mature domain theory, regardless of the model architecture employed. Theoretical understanding of protein physics, atmospheric dynamics, and molecular chemistry is what enables the embedding of constraints during training, the calibration of uncertainty metrics, and the interpretation of systematic failure patterns. Moreover, this theoretical knowledge enters the reliability-supporting workflow at multiple stages, not only during model training but also in filtering outputs, designing validation experiments, and interpreting results. Where such theory is available, as in structural biology and meteorology, generative AI can support reliable inference. Where it is not, alternative approaches will be needed.
@@
Second, about the models themselves: A closely related question is whether the models must be domain-specific, or whether general-purpose architectures could eventually achieve similar reliability when embedded in appropriate workflows. On this question I remain agnostic. While domain-specific models currently outperform general-purpose models (largely because domain-specific architectures naturally align with domain-specific workflows), there are reasons to take seriously the possibility that general-purpose models might eventually achieve comparable reliability. Multi-modal models such as GPT-4 and Gemini 2.5 exhibit capacities that cannot be replicated by chaining together narrow tools. Their integration of language, image, and video appears to produce synergies within the model’s latent space that surpass what modular composition can deliver. Whether chemical or physical representations could be treated as additional modalities in this sense, and whether such integration, when combined with domain-specific theoretical scaffolding at the workflow level, would yield genuinely new epistemic benefits, remains an open question worth exploring.
@@
The central argument of this paper reflects the Deweyan insight quoted in the epigraph: scientific knowledge is not a matter of grasping antecedently given certainties, but of developing ways to make sure—methods for identifying error and managing uncertainty. Generative AI calls for new methods of doing that work, and simultaneously makes that work more difficult. [^12] This dependency on foundational science suggests a cautionary note about research funding. While enthusiasm for AI-driven science grows, there is a risk of reallocating funding away from traditional theory-building and experimentation. Yet generative AI models achieve reliability precisely because they are embedded in theoretical frameworks developed through decades of foundational research. The epistemic scaffolding that AI-based inference depends on requires continued investment. Neglecting these foundations would be self-defeating. Nevertheless, as AlphaFold and GenCast demonstrate, there is reason to hope that we will be equal to the task.